# A Quadratic Programming Approach to Flight Envelope Protection Using Control Barrier Functions

### v2

@@ Abstract
Ensuring the safe operation of aerospace systems within their prescribed flight envelope is a fundamental requirement for modern flight control systems. Flight envelope protection prevents violations of aerodynamic, structural, and performance constraints, mitigating risks such as stall, excessive loads, and loss of control. Conventional FEP approaches, such as reference clipping via saturation functions and model-based command filtering, impose constraints at the reference input level but often fail to account for closed-loop system dynamics, potentially leading to constraint violations during transients. This paper introduces a new approach to the flight envelope protection problem by employing a quadratic programming-based safety filter using control barrier functions to dynamically enforce flight envelope constraints while preserving control performance. Unlike traditional reference filtering methods, the control barrier function-based safety filter actively ensures strict forward invariance of the safe flight envelope set, integrating seamlessly with existing control architectures. The proposed framework is implemented in a nonlinear missile flight control system and evaluated in a simulated environment. The results demonstrate its ability to prevent constraint violations while minimizing conservatism, offering a robust alternative to existing flight envelope protection methodologies.
@@ 1 Introduction
Flight envelope protection (FEP) plays a crucial role in this context, ensuring that aerospace vehicles remain within safe operational limits defined by aerodynamic, structural, and performance constraints. The flight envelope encompasses critical parameters, including speed, load factor, altitude, and angle of attack, all of which must be respected to prevent catastrophic system failure, such as structural damage, aerodynamic stall, or loss of control [1, 2]. Given the increasing complexity of modern aerospace platforms, robust and theoretically grounded methods for FEP are crucial to ensure operational safety while enabling vehicles to fully realize their performance potential. In most cases, FEP has been implemented at the attitude control level using reference filtering techniques [3, 4]. Two widely used approaches in this context are reference clipping (RC) via saturation functions and model-based command filtering (MBCF). The first approach, RC, applies saturation functions that limit the reference commands based on predefined upper and lower bounds. This method ensures that commanded values do not exceed safe limits, effectively enforcing a constraint at the reference input level [5]. While straightforward to implement, the RC approach fundamentally transforms the FEP problem into a reference tracking problem, leading to several limitations. Most notably, it does not account for the closed-loop system dynamics, which can result in constraint violations during transient phases, as observed in [2, 6]. Additionally, since the clipping function operates independently of the system dynamics, it lacks theoretical guarantees of strict constraint satisfaction under disturbances and unmodeled dynamics. An alternative method is MBCF, such as command filters or model reference filters [7]. These approaches incorporate an assumed dynamic model of the vehicle to shape the reference commands, enabling smoother control actions that respect dynamic constraints. When properly designed, such filters can provide empirical and theoretical guarantees of forward invariance, ensuring that the system remains within safe operating limits, as shown in [8], [9] and [10]. However, the complexity of designing these filters increases significantly, particularly for systems with relative degrees greater than two.
@@
Recent advancements in control theory offer an alternative perspective on designing FEP systems, framing the problem as a closed-loop state-constrained control problem. In such a framework, the controllers not only track desired set points but also dynamically enforce safety constraints throughout the system’s operation, both during steady-state and transient phases. Control Barrier Functions (CBFs) have emerged as a powerful tool within this class of safety-critical control methods. CBFs guarantee the forward invariance of a defined safe set, providing formal assurances that the system remains within prescribed limits under all operational conditions [11]. The optimization-based nature of CBFs allows for seamless integration of a wide range of constraints, including physical limitations such as load factor and stall prevention, as well as operational constraints like altitude limits and obstacle avoidance [12, 13]. A significant advantage of CBF-based methods is their ability to augment existing control architectures. Rather than requiring a complete redesign, CBFs can function as an additional safety layer, leveraging the existing closed-loop dynamics as part of the optimization problem [14, 15]. This compatibility facilitates their implementation in modern aerospace systems, while the real-time adaptability of CBFs enables optimal performance under varying conditions, disturbances, and uncertainties [16, 11, 17]. Despite their potential, the application of CBFs to FEP remains relatively underexplored in the flight control community.
@@ 2.1 The Mathematical Concept of Control Barrier Functions
We refer to $S$ as the safe set of the considered dynamics, given in Eq. (1). The following definitions are introduced [18, 19]:
@@
[…]
We use Definition 4 to introduce Nagumo’s theorem[18]:
@@
where $g:\chi\to\mathbb{R}^{n\times m}$ is Lipschitz and $u(t)\in\mathbb{R}^{m}$. We introduce the notion of a CBF such that its existence allows the system to be rendered safe w.r.t. $S$ [11, 14] and allows a weaker requirement for system safety with a control input $u(t)$, similar to (8).
@@
$\dot{h}(x(t),u(t))=\frac{\partial h}{\partial x}\begin{bmatrix}f(x(t))+g(x(t))u(t)\end{bmatrix}=L_{f}h(x(t))+L_{g}h(x(t))u(t)\vspace{0.2cm}$ (11)
As discussed in [20, 21], any Lipschitz feedback controller in the form of:
@@
renders the system forward invariant within $S$. One way to obtain such a controller is to use the universal input-to-state safeguarding formula, as proposed in [21], which delivers the following control law:
@@
Since the relevant nonlinear dynamics of the flight systems considered in this work are of a relative degree greater than one, we extend the discussion of CBFs with the notion of High Order Barrier Functions (HOBFs) and High Order Control Barrier Functions (HOCBFs), as proposed in [22]. The control-theoretical construct of CBFs introduced here, along with their extension to HOCBFs, provide a unified methodology for enforcing state constraints derived from operational limits, such as aerodynamic, structural, and performance constraints, which are critical for ensuring safe flight with complex and nonlinear behaviors. By establishing these theoretical foundations, the framework sets the stage for its application to practical systems, as demonstrated in this work through its integration into an FEP scheme. we consider a time-varying function to define an invariant set for system (1). For a $d$-th order differentiable function $h:\mathbb{R}^{n}\times[t_{0},\infty)\to\mathbb{R}$ (where $t_{0}$ denotes the initial time), we define a series of functions $\psi_{0},\psi_{1},\ldots,\psi_{d}$ recursively as:
@@
for all $(x(t),t)\in S_{1}(t)\cap S_{2}(t)\cap\cdots\cap S_{d}(t)\times[t_{0},\infty)$. In the above equation, $O(h(x(t),t))$ denotes the remaining Lie derivatives along $f$ and partial derivatives with respect to $t$ with degree less than or equal to $d-1$.
@@ 2.2 Considered Nonlinear Missile Dynamics
This study considers an axially symmetric skid-to-turn tail-controlled missile system, as illustrated in Fig. 2   [23].
@@
Under the assumption that the roll motion is rapidly stabilized, the missile motions can be decoupled into two perpendicular and identical channels (i.e., pitch and yaw). The missile pitch dynamics model utilized in this study is an adjusted version of the dynamical model shared in [24]. The model is a reduced-order version of a six-degree-of-freedom model, where non-relevant dynamics have been removed.
@@
where $\alpha(t)$ and $q(t)$ represent the angle of attack and the pitch rate, which are the state variables. Furthermore, $\delta(t)$ represents the control fin deflection, which is the control input of the missile system, and $n_{z}$ denotes the lateral load. V represent the missile velocity $Q$ represents the dynamic pressure ($\rho V^{2}/2$). The aerodynamic coefficients $C_{Z,\alpha}$, $C_{Z,\delta}$,$C_{M,\alpha}$, $C_{M,q}$, and $C_{M,\delta}$ are obtained from wind-tunnel testing or aero-prediction methods. $S$, $d$, $m$, and $I_{yy}$ refer to the constant reference area, diameter, mass, and moment of inertia, respectively.
@@
$C_{Z,\alpha}$32.5925
$C_{Z,\delta}$7.1863
$C_{m,0}$3.277
$C_{m,\alpha}$80.4716
$C_{m,q}$56.1499
$C_{m,\delta}$69.6272
@@ 4.1 Nonlinear Attitude Controller Design
For the sake of reproducibility of the results presented here, it was decided to control the nonlinear longitudinal dynamics of the missile using a simple cascaded nonlinear dynamic inversion (NDI) control strategy, as presented in [25]. However, it is worth noting that the choice of attitude controller can be viewed as separate from the FEP problem, and any performance-oriented attitude controller could be integrated into the proposed overall flight control architecture.
The fundamental principle of NDI-based control systems is to transform the nonlinear input-output relationship into a linear relationship through state feedback and coordinate transformation. This method contrasts with the traditional Jacobian linearization, which relies on local linear approximations at specific operating points within the flight envelope [26]. By leveraging nonlinear control techniques, the requirement for gain scheduling is eliminated, resulting in a simplified system response with controlled variables exhibiting integrator-like dynamics [27]. The cascaded controller architecture is depicted in Fig. 5.
@@
As discussed in Section 1, two widely used methods in classical FEP are RC via saturation functions and MBCF. While these two methods aim to enforce constraints by shaping the commanded trajectories, their effectiveness depends critically on the dynamic behavior of the closed-loop system. In particular, tracking delays, steady-state errors, or overshoots in response to reference commands can result in constraint violations—even when the reference signal itself remains within nominally safe bounds. To explicitly analyze this dependency later on in this work, two distinct gain configurations of the previously introduced cascaded NDI controller are considered. The controller gains have been defined using a standard Linear Quadratic Regulator (LQR) technique [28]. The first gain configuration is referred to as the "conservative" control setting, characterized by lower gain values. While this tuning results in a conservative and less aggressive control response, it leads to slow reference tracking and smaller overshoots in tracking. The second configuration is labeled the "aggressive" control setting, which uses higher gain values to yield faster tracking performance. In this case, the system response is considerably more agile, closely following the reference signal with minimal phase lags, albeit at the cost of a transient overshoot. The relevant parameters for the considered controller settings are given in Table 2.
@@ 4.2 CBF-Based Nonlinear Safety Filter Design
In the integrated QP-based safety filter formulation, the upper and lower bounds on $\alpha(t)$ are enforced via two distinct HOCBF-derived inequality constraints. In general, this setup could lead to feasibility issues in the QP, especially when both constraints become simultaneously active or conflicting. However, due to the structure of the considered FEP problem, this conflict is typically avoided in practice. Specifically, the system either approaches the upper or the lower bound of the operational $\alpha(t)$ limits depending on the direction of motion—i.e., whether $\alpha(t)$ is increasing or decreasing. As a result, only one of the two HOCBF constraints is usually active at a given time, which significantly reduces the risk of mutual constraint conflict. Nevertheless, additional care must be taken to ensure feasibility when input constraints are present. As discussed in [29], feasibility guarantees can be ensured through careful selection of the design parameters for the integrated CBFs. This is further discussed in the following part of this section, which outlines a tuning strategy to maintain forward invariance while preserving the feasibility of the integrated safety filter.
@@ 4.2.1 Derivation of CBF-based Safety Constraints
Under consideration of aerodynamic coefficients, which primarily depend on the Mach number, for which we can also assume slowly changing properties relative to the considered attitude dynamics, we assume [30]:
@@
As discussed in [22, 31], the construction of valid HOCBFs for systems with relative degree greater than one and limited control authority requires careful tuning of the associated design parameters. The choice of these parameters is essential to ensure that the resulting inequality constraints yield forward-invariant safe sets under actuation constraints.
@@
[…]

### v3

# Quadratic Programming Approach to Flight Envelope Protection Using Control Barrier Functions

@@ Abstract
Ensuring the safe operation of aerospace systems within their prescribed flight envelope is a fundamental requirement for modern flight control systems. Flight envelope protection (FEP) prevents violations of aerodynamic, structural, and performance constraints, mitigating risks such as stall, excessive loads, and loss of control. Conventional FEP approaches, such as reference clipping via saturation functions and model-based command filtering, impose constraints at the reference input level but often fail to account for closed-loop system dynamics, potentially leading to constraint violations during transients. This paper introduces a new approach to flight envelope protection by employing a quadratic-programming-based safety filter using control barrier functions to dynamically enforce flight envelope constraints while preserving control performance. Unlike traditional reference filtering methods, the proposed control barrier function-based safety filter actively ensures forward invariance of the safe flight envelope set while seamlessly integrating with existing control architectures. The framework is implemented in a nonlinear missile flight control system and evaluated in a simulated environment. The results demonstrate its ability to prevent constraint violations while minimizing conservatism, offering a robust alternative to existing flight envelope protection methodologies.
@@ 1 Introduction
Flight envelope protection (FEP) plays a crucial role in this context, ensuring that aerospace vehicles remain within safe operational limits defined by aerodynamic, structural, and performance constraints. The flight envelope encompasses critical parameters, including speed, load factor, altitude, and angle of attack, all of which must be respected to prevent catastrophic system failure, such as structural damage, aerodynamic stall, or loss of control [Oudin2017, Stougie2024-ym]. Given the increasing complexity of modern aerospace platforms, robust and theoretically grounded methods for FEP are crucial to ensure operational safety while enabling vehicles to fully realize their performance potential. In most cases, FEP has been implemented at the attitude control level using reference filtering techniques [Tang2009, Falkena2010]. Two widely used approaches in this context are reference clipping (RC) via saturation functions and model-based command filtering (MBCF). The first approach, RC, applies saturation functions that limit the reference commands based on predefined upper and lower bounds. This method ensures that commanded values do not exceed safe limits, effectively enforcing a constraint at the reference input level [Seo2017]. While straightforward to implement, the RC approach fundamentally transforms the FEP problem into a reference tracking problem, leading to several limitations. Most notably, it does not account for the closed-loop system dynamics, which can result in constraint violations during transient phases, as observed in [Stougie2024-ym, Steffensen2019]. Additionally, since the clipping function operates independently of the system dynamics, it lacks theoretical guarantees of strict constraint satisfaction under disturbances and unmodeled dynamics. An alternative method is MBCF, such as command filters or model reference filters [Lombaerts2017]. These approaches incorporate an assumed dynamic model of the vehicle to shape the reference commands, enabling smoother control actions that respect dynamic constraints. When properly designed, such filters can provide empirical and theoretical guarantees of forward invariance, ensuring that the system remains within safe operating limits, as shown in [Grondman2018], [Autenrieb_2024b] and [Autenrieb2024]. However, the complexity of designing these filters increases significantly, particularly for systems with relative degrees greater than two.
@@
Recent advancements in control theory offer an alternative perspective on designing FEP systems, framing the problem as a closed-loop state-constrained control problem. In such a framework, the controllers not only track desired set points but also dynamically enforce safety constraints throughout the system’s operation, both during steady-state and transient phases. Control Barrier Functions (CBFs) have emerged as a powerful tool within this class of safety-critical control methods. CBFs guarantee the forward invariance of a defined safe set, providing formal assurances that the system remains within prescribed limits under all operational conditions [Ames_2014]. The optimization-based nature of CBFs allows for seamless integration of a wide range of constraints, including physical limitations such as load factor and stall prevention, as well as operational constraints like altitude limits and obstacle avoidance [Agrawal2021-aw, autenrieb2025_arxiv]. A significant advantage of CBF-based methods is their ability to augment existing control architectures. Rather than requiring a complete redesign, CBFs can function as an additional safety layer, leveraging the existing closed-loop dynamics as part of the optimization problem [Ames_2017, Li2023-wy]. This compatibility facilitates their implementation in modern aerospace systems, while the real-time adaptability of CBFs enables optimal performance under varying conditions, disturbances, and uncertainties [Taylor2020-em, Ames_2014, solanocastellanos2024]. Despite their potential, the application of CBFs to FEP remains relatively underexplored in the flight control community.
@@ 2.1 The Mathematical Concept of Control Barrier Functions
We refer to $S$ as the safe set of the considered dynamics, given in Eq. (1). The following definitions are introduced [Nagumo_1942, Blanchini_1999]:
@@
If $S$ is convex $T_{S}(x(t))$ is convex, and $\liminf$” can be replaced by $\lim$”. Furthermore if $x\in int(S)$, then $T_{S}(x(t))=\mathbb{R}^{n}$, whereas if $x\notin S$, then $T_{S}(x(t))=\emptyset$, since $S$ is defined as a closed set. Therefore, $T_{S}(x(t))$ is only non-trivial on the boundary of $S$.
We use Definition 4 to introduce Nagumo’s theorem[Nagumo_1942]:
@@
where $g:\chi\to\mathbb{R}^{n\times m}$ is Lipschitz and $u(t)\in\mathbb{R}^{m}$. We introduce the notion of a CBF such that its existence allows the system to be rendered safe w.r.t. $S$ [Ames_2014, Ames_2017] and allows a weaker requirement for system safety with a control input $u(t)$, similar to (8).
@@
$\dot{h}(x(t),u(t))=\frac{\partial h}{\partial x}\begin{bmatrix}f(x(t))+g(x(t))u(t)\end{bmatrix}=L_{f}h(x(t))+L_{g}h(x(t))u(t)\vskip 5.69046pt$ (11)
As discussed in [XU2015, Kolathaya2019], any Lipschitz feedback controller in the form of:
@@
renders the system forward invariant within $S$. One way to obtain such a controller is to use the universal input-to-state safeguarding formula, as proposed in [Kolathaya2019], which delivers the following control law:
@@
with $A(x(t))=L_{f}h(x(t))+\alpha(h(x(t)))$ and $B(x(t))=L_{g}h(x(t))^{T}$, which leads to $\dot{h}(x(t),u(t))\geq 0$ at the boundaries. Hence, the system always stays within the safe set $S$. Alternatively, the authors in [Ames_2014] proposed a real-time optimization-based safety filter to balance control performance and safety optimally. The formulated CBF-based quadratic program (QP-CBF) problem minimizes the Euclidean norm between a desired control input u(t)u*(t) (for example, coming from a stable and performance-oriented nonlinear controller) and the safe control input $u(t)$:
@@
[…]
$A_{u}u(t)+b\leq 0,$
with $A_{u}u(t)+b$ representing box constraints on the control input $u$ to address the problem of a limited admissible set.
Since the relevant nonlinear dynamics of the flight systems considered in this work are of a relative degree greater than one, we extend the discussion of CBFs with the notion of High Order Barrier Functions (HOBFs) and High Order Control Barrier Functions (HOCBFs), as proposed in [Xiao_2022hocbf]. The control-theoretical construct of CBFs introduced here, along with their extension to HOCBFs, provide a unified methodology for enforcing state constraints derived from operational limits, such as aerodynamic, structural, and performance constraints, which are critical for ensuring safe flight with complex and nonlinear behaviors. By establishing these theoretical foundations, the framework sets the stage for its application to practical systems, as demonstrated in this work through its integration into an FEP scheme. we consider a time-varying function to define an invariant set for system (1). For a $d$-th order differentiable function $h:\mathbb{R}^{n}\times[t_{0},\infty)\to\mathbb{R}$ (where $t_{0}$ denotes the initial time), we define a series of functions $\psi_{0},\psi_{1},\ldots,\psi_{d}$ recursively as:
@@
for all $(x(t),t)\in S_{1}(t)\cap S_{2}(t)\cap\cdots\cap S_{d}(t)\times[t_{0},\infty)$. In the above equation, $O(h(x(t),t))$ denotes the remaining Lie derivatives along $f$ and partial derivatives with respect to $t$ with degree less than or equal to d1d-1.
@@ 2.2 Considered Nonlinear Missile Dynamics
This study considers an axially symmetric skid-to-turn tail-controlled missile system, as illustrated in Fig. 2 [Lee_2016].
@@
Under the assumption that the roll motion is rapidly stabilized, the missile motions can be decoupled into two perpendicular and identical channels (i.e., pitch and yaw). The missile pitch dynamics model utilized in this study is an adjusted version of the dynamical model shared in [Hwang_2017]. The model is a reduced-order version of a six-degree-of-freedom model, where non-relevant dynamics have been removed.
@@
where $\alpha(t)$ and $q(t)$ represent the angle of attack and the pitch rate, which are the state variables. Furthermore, $\delta(t)$ represents the control fin deflection, which is the control input of the missile system, and $n_{z}$ denotes the lateral load. V represent the missile velocity $Q$ represents the dynamic pressure ($\rho V^{2}/2$). The aerodynamic coefficients $C_{Z,\alpha}$, $C_{Z,\delta}$$C_{M,\alpha}$, $C_{M,q}$, and $C_{M,\delta}$ are obtained from wind-tunnel testing or aero-prediction methods. $S$, $d$, $m$, and $I_{yy}$ refer to the constant reference area, diameter, mass, and moment of inertia, respectively.
@@
$C_{Z,\alpha}$ -32.5925
$C_{Z,\delta}$ -7.1863
$C_{m,0}$ -3.277
$C_{m,\alpha}$ -80.4716
$C_{m,q}$ -56.1499
$C_{m,\delta}$ -69.6272
@@ 4.1 Nonlinear Attitude Controller Design
For the sake of reproducibility of the results presented here, it was decided to control the nonlinear longitudinal dynamics of the missile using a simple cascaded nonlinear dynamic inversion (NDI) control strategy, as presented in [daCosta2003]. However, it is worth noting that the choice of attitude controller can be viewed as separate from the FEP problem, and any performance-oriented attitude controller could be integrated into the proposed overall flight control architecture.
The fundamental principle of NDI-based control systems is to transform the nonlinear input-output relationship into a linear relationship through state feedback and coordinate transformation. This method contrasts with the traditional Jacobian linearization, which relies on local linear approximations at specific operating points within the flight envelope [Acquatella_2022]. By leveraging nonlinear control techniques, the requirement for gain scheduling is eliminated, resulting in a simplified system response with controlled variables exhibiting integrator-like dynamics [Holzapfel_2004]. The cascaded controller architecture is depicted in Fig. 5.
@@
As discussed in Section 1, two widely used methods in classical FEP are RC via saturation functions and MBCF. While these two methods aim to enforce constraints by shaping the commanded trajectories, their effectiveness depends critically on the dynamic behavior of the closed-loop system. In particular, tracking delays, steady-state errors, or overshoots in response to reference commands can result in constraint violations—even when the reference signal itself remains within nominally safe bounds. To explicitly analyze this dependency later on in this work, two distinct gain configurations of the previously introduced cascaded NDI controller are considered. The controller gains have been defined using a standard Linear Quadratic Regulator (LQR) technique [Kalman_1960LQR]. The first gain configuration is referred to as the "conservative" control setting, characterized by lower gain values. While this tuning results in a conservative and less aggressive control response, it leads to slow reference tracking and smaller overshoots in tracking. The second configuration is labeled the "aggressive" control setting, which uses higher gain values to yield faster tracking performance. In this case, the system response is considerably more agile, closely following the reference signal with minimal phase lags, albeit at the cost of a transient overshoot. The relevant parameters for the considered controller settings are given in Table 2.
@@ 4.2 CBF-Based Nonlinear Safety Filter Design
In the integrated QP-based safety filter formulation, the upper and lower bounds on $\alpha(t)$ are enforced via two distinct HOCBF-derived inequality constraints. In general, this setup could lead to feasibility issues in the QP, especially when both constraints become simultaneously active or conflicting. However, due to the structure of the considered FEP problem, this conflict is typically avoided in practice. Specifically, the system either approaches the upper or the lower bound of the operational $\alpha(t)$ limits depending on the direction of motion—i.e., whether $\alpha(t)$ is increasing or decreasing. As a result, only one of the two HOCBF constraints is usually active at a given time, which significantly reduces the risk of mutual constraint conflict. Nevertheless, additional care must be taken to ensure feasibility when input constraints are present. As discussed in [Spiller_2025], feasibility guarantees can be ensured through careful selection of the design parameters for the integrated CBFs. This is further discussed in the following part of this section, which outlines a tuning strategy to maintain forward invariance while preserving the feasibility of the integrated safety filter.
@@ 4.2.1 Derivation of CBF-based Safety Constraints
Under consideration of aerodynamic coefficients, which primarily depend on the Mach number, for which we can also assume slowly changing properties relative to the considered attitude dynamics, we assume [Lee2020]:
@@
As discussed in [Xiao_2022hocbf, Agrawal2021], the construction of valid HOCBFs for systems with relative degree greater than one and limited control authority requires careful tuning of the associated design parameters. The choice of these parameters is essential to ensure that the resulting inequality constraints yield forward-invariant safe sets under actuation constraints.
@@
However, as illustrated in the lower subplots, the different parameter configurations result in different demands on the available actuators. While magnitude constraints are not critical in this scenario, the actuator rate limits influence the feasibility of the safety filter. For instance, the parameter setting $(\gamma_{1},\gamma_{2})=(3,\ 5)$, shown in yellow, leads to a more aggressive safety enforcement, pushing the control rate $\dot{\delta}$ at one point to its limit. This behavior is further clarified in the magnified inset (bottom right), showing that the system operates at the edge of feasibility. While such a setting ensures minimal compromise on performance while rendering the system safe, it may not be desirable in practice due to excessive control activity. Therefore, the aim is to identify a setting that balances safety with minimal intervention and reduced actuation effort. In the regarded case, the setting $(\gamma_{1},\gamma_{2})=(2,\ 1)$ (orange curve) yields a favorable trade-off—preserving constraint satisfaction without saturating the rate limits, suggesting the further use of this parameter setting for the CBF-based safety filter.